# An example of the Jantzen filtration of a D-module

### v1

@@ 1.1. Structure of the paper
The focus of this paper is an extended $\operatorname{SL}_{2}(\mathbb{C})$ example, and the majority of the exposition is dedicated to the computation of the geometric Jantzen filtration in this example. We begin the paper in §2 with a brief outline of the construction of the geometric Jantzen filtration for a general Lie algebra. In §3, we shift our focus to $\operatorname{SL}_{2}(\mathbb{C})$ for the remainder of the paper. Our first computation in §$\S 3.1$ establishes an algebra homomorphism from the extended universal enveloping algebra to global differential operators on base affine space. This algebra homomorphism is what allows us to view the global sections of $\mathcal{D}$-modules as modules over the (extended) universal enveloping algebra. In §3.2, we give some background on $H$-monodromic $\mathcal{D}$-modules, and explain their relationship to modules over twisted sheaves of differential operators. In §3.3, we introduce the $\mathcal{D}$-modules whose global sections contain Verma modules and dual Verma modules — these are the $\mathcal{D}$-modules which we will endow with geometric Jantzen filtrations. We illustrate the $\mathfrak{sl}_{2}(\mathbb{C})$-module structure on their global sections in Figures 1 and 2. In §3.4, we introduce the maximal extension functor, which gives us the deformation necessary for the construction of the Jantzen filtration. We compute the maximal extension of the structure sheaf on an open subset of base affine space, and illustrate in Figures 3 and 4 how deformed Verma modules and deformed dual Verma modules arise in this computation. We illustrate in Figures 5 and 6 the global sections of the maximal extension, identifying them with the big projective in category $\mathcal{O}$. In §3.5, we introduce monodromy filtrations in abelian categories, and define the geometric Jantzen filtration on a $\mathcal{D}$-module using this set-up. We compute the monodromy filtration of the maximal extension, and illustrate its global sections in Figure 7. This specializes to the geometric Jantzen filtration on certain sub- and quotient sheaves. We finish in §3.6 by relating the geometric Jantzen filtration to the algebraic one. This step is critical, and is given only minimal justification in [BB93]. We introduce the algebraic Jantzen filtration on a Verma module in §3.6.1, then explain why the global sections of the geometric Jantzen filtration align with the algebraic Jantzen filtration in §3.6.2. Figure 8 illustrates this relationship in our example.
2. The Jantzen filtration of a $\mathcal{D}$-module
This paper is concerned with the example of $\operatorname{SL}_{2}(\mathbb{C})$. However, some amount of general theory will be helpful to set the scene. In this section, we briefly orient the reader with the necessary general theory. More details of this construction can be found in [BB93].
2.1. The monodromy filtration
Geometric Jantzen filtrations are intimately related to monodromy filtrations. Given an object $A$ in an abelian category $\mathcal{A}$ and a nilpotent endomorphism $s\in\operatorname{End}(A)$, there is a unique filtration of $A$ satisfying certain symmetry properties with respect to $s$. (See §3.5 for a precise definition.) This is the monodromy filtration of $A$. It induces filtrations on the sub and quotient objects $\ker(s:A\rightarrow A)$ and $\operatorname{coker}(s:A\rightarrow A)$ in the natural way.
2.2. Geometric Jantzen filtration
Certain $\mathcal{D}$-modules come naturally equipped with nilpotent endomorphisms, and thus acquire monodromy filtrations. If $Y$ is a smooth algebraic variety with a fixed regular function $f:Y\rightarrow\mathbb{A}^{1}$, the maximal extension $\Xi_{f}\mathcal{M}_{U}$ of a holonomic $\mathcal{D}_{U}$-module $\mathcal{M}_{U}$ on $U=f^{-1}(\mathbb{A}^{1}-\{0\})$ is an object in the abelian category of holonomic $\mathcal{D}_{Y}$-modules which has a natural nilpotent endomorphism $s$. Hence it has a monodromy filtration. Moreover, by the construction of the maximal extension functor (which is outlined in §3.4), $\ker(s:\Xi_{f}\mathcal{M}_{U}\rightarrow\Xi_{f}\mathcal{M}_{U})=j_{!}\mathcal{M}_{U}$ and $\operatorname{coker}(s:\Xi_{f}\mathcal{M}_{U}\rightarrow\Xi_{f}\mathcal{M}_{U})=j_{+}\mathcal{M}_{U}$, where $j:U\hookrightarrow Y$ is inclusion. In this way, we obtain filtrations of the $\mathcal{D}_{Y}$-modules $j_{!}\mathcal{M}_{U}$ and $j_{+}\mathcal{M}_{U}$ from the monodromy filtration of $\Xi_{f}\mathcal{M}_{U}$. These are the geometric Jantzen filtrations.
2.3. Geometric Jantzen filtration on Harish-Chandra sheaves
If $(\mathfrak{g},K)$ is a Harish-Chandra pair [^1] Meaning that $\mathfrak{g}$ is a complex semisimple Lie algebra and $K$ is an algebraic group equipped with a morphism $K\rightarrow\mathrm{Int}\mathfrak{g}$ whose differential is injective, with the property that $K$ acts with finitely many orbits on the flag variety of $\mathfrak{g}$. with admissible $K$-orbits [^2] See [BB93, 3.4.1(i)] for this rather technical definition. In the cases where $K=N$ is a maximal nilpotent subgroup or $(\mathfrak{g},K)$ is a symmetric pair, all orbits are admissible. in the flag variety $X$ of $\mathfrak{g}$, then the !!!- and $+$-standard $\mathcal{D}_{X}$-modules obtained by pushing forward $K$-equivariant connections on $K$-orbits can be made to fit into the framework of §2.2, and thus acquire geometric Jantzen filtrations. To explain more details of this construction we must introduce more notation.
Let $G$ be the simply connected semisimple Lie group associated to $\mathfrak{g}$, and fix a Borel subgroup $B\subset G$. Let $N\subset B$ denote the unipotent radical of $B$, and set $H:=B/N$ to be the abstract maximal torus of $G$ [CG97, Lem. 6.1.1], with Lie algebra $\mathfrak{h}:=\operatorname{Lie}H$. Let $\widetilde{X}:=G/N$ be the base affine space corresponding to $\mathfrak{g}$, and identify $X$ with $G/B$. Let $\pi:\widetilde{X}\rightarrow X$ be projection. This is a principal $G$-equivariant $H$-bundle with respect to the right action of $H$ on $\widetilde{X}$ by right multiplication.
For a $K$-orbit $Q$ in $X$, denote by $\widetilde{Q}=\pi^{-1}(Q)$ the corresponding union of $K$-orbits in $\widetilde{X}$. A choice of an integral weight $\lambda\in\mathfrak{h}^{*}$ (the “deformation direction”) determines a function $f_{\lambda}:\overline{\widetilde{Q}}\rightarrow\mathbb{A}^{1}$ on the closure of $\widetilde{Q}$ such that $f_{\lambda}^{-1}(\mathbb{A}^{1}-\{0\})=\widetilde{Q}$. This function extends to a regular function on $\widetilde{X}$, which determines a maximal extension functor $\Xi_{f_{\lambda}}:\mathcal{M}_{\mathrm{hol}}(\mathcal{D}_{U})\rightarrow\mathcal{M}_{\mathrm{hol}}(\mathcal{D}_{\widetilde{X}})$, where $U$ is the preimage in $\widetilde{X}$ of $\mathbb{A}^{1}-\{0\}$ under the extension of $f_{\lambda}$, as in §2.2. Restricting $\Xi_{f_{\lambda}}$ to the category of holonomic $\mathcal{D}_{U}$-modules supported on $\widetilde{Q}$ results in a functor
$\Xi_{f_{\lambda}}:\mathcal{M}_{\mathrm{hol}}(\mathcal{D}_{\widetilde{Q}})\rightarrow\mathcal{M}_{\mathrm{hol}}(\mathcal{D}_{\overline{\widetilde{Q}}}).$
Let $\mathcal{M}_{\widetilde{Q}}$ be an irreducible $K$-equivariant connection on $\widetilde{Q}$. Using the procedure outlined above, the modules $j_{\widetilde{Q}!}\mathcal{M}_{\widetilde{Q}}$ and $j_{\widetilde{Q}+}\mathcal{M}_{\widetilde{Q}}$ acquire from $\Xi_{f_{\lambda}}\mathcal{M}_{\widetilde{Q}}$ geometric Jantzen filtrations. Here $j_{\widetilde{Q}}:\widetilde{Q}\hookrightarrow\overline{\widetilde{Q}}$ is inclusion of $\widetilde{Q}$ into its closure. Because $\overline{\widetilde{Q}}$ is closed in $\widetilde{X}$, we can use Kashiwara’s theorem to lift these filtrations to filtrations of the standard $K$-equivariant $\mathcal{D}_{\widetilde{X}}$-modules $i_{\widetilde{Q}!}\mathcal{M}_{\widetilde{Q}}$ and $i_{\widetilde{Q}+}\mathcal{M}_{\widetilde{Q}}$, for $i_{\widetilde{Q}}:\widetilde{Q}\hookrightarrow\widetilde{X}$ the inclusion.
2.4. Relationship to algebraic Jantzen filtration
Classical results in Lie theory [Bei81, Milb] relate global sections of standard $K$-equivariant $\mathcal{D}_{X}$-modules to well-known classes of $\mathfrak{g}$-modules, such as Verma modules or principal series representations, depending on specific choices of $K$. These $\mathfrak{g}$-modules also admit algebraic Jantzen filtrations (see §3.6.1), which are usually defined in terms of contravariant forms. Under Beilinson–Bernstein localisation, the algebraic and geometric Jantzen filtrations align. In §3.6.2 we explain this relationship in our example.
@@ 3. Example
  • If $Y$ is normal with $\mathrm{codim}_{\overline{Y}}(\overline{Y}\backslash Y)\geq 2$, then $\mathcal{O}(Y)=\mathcal{O}(\overline{Y})$ and $\mathcal{D}(Y)=\mathcal{D}(\overline{Y})$. (In particular, for $Y=\widetilde{X}$, this implies that global differential operators are nothing more than the Weyl algebra in 2 variables [^3] Outside of $\mathfrak{sl}_{2}(\mathbb{C})$, the situation is less straightforward. For a Lie algebra $\mathfrak{g}\neq\mathfrak{sl}_{2}(\mathbb{C})^{m}$, the affine closure of the corresponding base affine space is singular, and the ring of global differential operators can be quite complicated, see, for example, [LS06]..)
  • The variety $\overline{Y}$ is $D$-affine, meaning that the global sections functor induces an equivalence of categories between the category of quasi-coherent $\mathcal{D}_{\overline{Y}}$-modules and the category of modules over $\mathcal{D}(\overline{Y})$.
  • Since the inclusion $i:Y\rightarrow\overline{Y}$ is an open immersion, the restriction functor $i^{+}$ on the corresponding categories of $\mathcal{D}$-modules is exact, and commutes with pushforwards from open affine subvarieties.
@@
Our strategy for gaining intuition about the $\mathcal{D}_{\widetilde{X}}$-modules arising in the construction of the Jantzen filtration is to illustrate the $\mathfrak{g}$-module structure on their global sections. This will give us an algebraic snapshot as to what is happening at each step in the construction sketched in Section 2. The first step is to differentiate the actions (3.7) and (3.8) to obtain a map $\mathcal{U}(\mathfrak{g})\otimes_{\mathcal{Z}(\mathfrak{g})}\mathcal{U}(\mathfrak{h})\rightarrow\Gamma(\widetilde{X},\mathcal{D}_{\widetilde{X}})$. This map provides the $\mathfrak{g}$-module structure on the global sections of $\mathcal{D}_{\widetilde{X}}$-modules. We dedicate this section to the computation of this map.
@@ 3.2. Monodromic $\mathcal{D}_{X}$-modules
An $H$-monodromic $\mathcal{D}_{X}$-module is a weakly $H$-equivariant [^4] For a reference on weakly equivariant $\mathcal{D}$-modules, see [MP98, §4]. $\mathcal{D}_{\widetilde{X}}$-module.
There is an equivalent characterization of monodromic $\mathcal{D}_{X}$-modules in terms of $H$-invariant differential operators which is established in [BB93, §2.5.2]. This perspective makes the structures of our examples more transparent, so we will take this approach to monodromicity. Below we describe the construction for $\mathfrak{g}=\mathfrak{sl}_{2}(\mathbb{C})$.
@@ 3.3. Verma modules and dual Verma modules
Using the map (3.18) constructed in Section 3.1, we can describe the $\widetilde{\mathcal{U}}$-module structure on various classes of $\mathcal{D}_{\widetilde{X}}$-modules. We will start by examining the $\mathcal{D}_{\widetilde{X}}$-modules $j_{+}\mathcal{O}_{U}$ and $j_{!}\mathcal{O}_{U}$, where $j:U\hookrightarrow\widetilde{X}$ is inclusion of the open $B$-orbit
@@
Here the $+$ and !!! indicate the $\mathcal{D}$-module push-forward functors, see [Mila]. These are the $\mathcal{D}_{\widetilde{X}}$-modules which will eventually be endowed with geometric Jantzen filtrations in Section 3.5. In this section, we describe the $\widetilde{\mathcal{U}}$-module structure on $\Gamma(\widetilde{X},j_{+}\mathcal{O}_{U})$ and $\Gamma(\widetilde{X},j_{!}\mathcal{O}_{U})$.
@@
This is a well-defined functor from the category of holonomic $\mathcal{D}_{Y}$-modules to itself [HTT08, Cor. 2.6.8].
@@
    The sheaf $\pi_{*}j_{!}\mathcal{O}_{U}$ is a monodromic $\mathcal{D}_{X}$-module. For each positive integer $n$, $\pi_{*}j_{!}\mathcal{O}_{U}$ has a subsheaf $(\pi_{*}j_{!}\mathcal{O}_{U})_{n}$ on which $R_{h}$ acts locally by the eigenvalue $n$. These subsheaves are modules over the TDO $\mathcal{D}_{n}$ (Remark 3.4). These are exactly the $\mathcal{D}_{n}$-modules appearing in [Rom21, §6, Fig. 3].
Remark 3.8.
It is interesting to contrast the computations above to the computations in [Rom21]. By working over base affine space instead of the flag variety, we are able to study all integral Verma and dual Verma modules at once using $j_{!}\mathcal{O}_{U}$ and $j_{+}\mathcal{O}_{U}$. This is different from the approach in [Rom21, §6], where Verma and dual Verma modules are studied one at a time by examining modules over the TDO $\mathcal{D}_{n}$ as $n$ varies. The relationship between these two approaches comes from Remark 3.4: one can view $\widetilde{\mathcal{D}}$ as a family of TDOs parametrized by $\mathfrak{h}^{*}$, and modules over $\widetilde{\mathcal{D}}$ can be viewed as families of modules over those TDOs.
@@ 3.4. The maximal extension $\Xi_{\rho}\mathcal{O}_{U}$
This functor (defined in (3.55)) extends $j_{+}$ and $j_{!}$ (see (3.57) - (3.58)), so it is a natural way to study both modules $j_{!}\mathcal{O}_{U}$ and $j_{+}\mathcal{O}_{U}$ at once. In this section, we give the construction of $\Xi_{\rho}$, then describe the $\widetilde{\mathcal{U}}$-module structure on $\Gamma(\widetilde{X},\Xi_{\rho}\mathcal{O}_{U})$.
@@
to be the $\mathcal{D}_{U}$-module obtained by twisting $\mathcal{M}_{U}$ by $I^{(n)}$. Note that $f^{s}\mathcal{M}_{U}^{(1)}=\mathcal{M}_{U}$.
@@
Now, we apply this general construction in the setting of our example. Let $\widetilde{X}$ and $U$ be as above (see (3.5) and (3.34)), and let $f_{\rho}$ be the function [^6] This choice of function corresponds to the deformation direction $\rho\in\mathfrak{h}^{*}$, see Remark 3.10..
@@ 3.4.2. Step 2: $+$-pushforward
annihilates any monomial basis element. (Here $\gamma_{\mathrm{HC}}$ is the Harish-Chandra projection in (3.3).) Hence, the module $\Gamma(\widetilde{X},j_{+}f^{s}\mathcal{O}_{U}^{(n)})$ has generalised, but not strict, infinitesimal character.

### v2

@@ 1.5. Structure of the paper
The remainder of the paper is dedicated to the computation of the geometric Jantzen filtration for the Lie algebra $\mathfrak{sl}_{2}(\mathbb{C})$. The computation is structured as follows.
§2.1: We establish an algebra homomorphism from the extended universal enveloping algebra to global differential operators on base affine space. This algebra homomorphism is what allows us to view the global sections of $\mathcal{D}_{\widetilde{X}}$-modules as modules over the (extended) universal enveloping algebra.
§2.2: We give some background on $H$-monodromic $\mathcal{D}_{X}$-modules, and explain their relationship to modules over twisted sheaves of differential operators.
§2.3: We introduce the $\mathcal{D}_{\widetilde{X}}$-modules whose global sections contain Verma modules and dual Verma modules — these are the $\mathcal{D}_{\widetilde{X}}$-modules which we will endow with geometric Jantzen filtrations. We illustrate the $\mathfrak{sl}_{2}(\mathbb{C})$-module structure on their global sections in Figures 1 and 2.
§2.4: We introduce the maximal extension functor, which gives the deformation necessary for the Jantzen filtration. We compute the maximal extension of the structure sheaf on an open subset of $\widetilde{X}$, and illustrate in Figures 3 and 4 how deformed Verma modules and deformed dual Verma modules arise geometrically. We illustrate in Figures 5 and 6 the global sections of the maximal extension, identifying them with the big projective in category $\mathcal{O}$.
§2.5: We define the geometric Jantzen filtration using monodromy filtrations. We compute the monodromy filtration of the maximal extension, and illustrate its global sections in Figure 7. This specialises to the geometric Jantzen filtration on certain sub- and quotient sheaves.
§2.6: We introduce the algebraic Jantzen filtration on a Verma module in §2.6.1, then explain why the global sections of the geometric Jantzen filtration align with the algebraic Jantzen filtration in §2.6.2. Figure 8 illustrates this relationship in our example.
1.6. Acknowledgements
This paper arose from computations in the first author’s honours thesis at the University of New South Wales. We would like to thank the referee for helpful suggestions which greatly improved the readability of the paper. The first author would like to thank Daniel Chan, who influenced his interests in this topic, and broadened his understanding of the various geometric techniques used in this paper. The second author would like to thank Jens Eberhardt, Adam Brown, and Geordie Williamson for many hours of conversations about Jantzen filtrations which contributed significantly to her understanding.
@@ 2. Example
  • If $Y$ is normal with $\mathrm{codim}_{\overline{Y}}(\overline{Y}\backslash Y)\geq 2$, then $\mathcal{O}(Y)=\mathcal{O}(\overline{Y})$ and $\mathcal{D}(Y)=\mathcal{D}(\overline{Y})$ [LS06, §2]. (In particular, for $Y=\widetilde{X}$, this implies that global differential operators are nothing more than the Weyl algebra in 2 variables [^6] Outside of $\mathfrak{sl}_{2}(\mathbb{C})$, the situation is less straightforward. For a Lie algebra $\mathfrak{g}\neq\mathfrak{sl}_{2}(\mathbb{C})^{m}$, the affine closure of the corresponding base affine space is singular, and the ring of global differential operators can be quite complicated, see, for example, [LS06]..)
  • Because the variety $\overline{Y}$ is affine, it is also $D$-affine, meaning that the global sections functor induces an equivalence of categories between the category of quasi-coherent $\mathcal{D}_{\overline{Y}}$-modules and the category of modules over $\mathcal{D}(\overline{Y})$.
  • Since the inclusion $i:Y\rightarrow\overline{Y}$ is an open immersion, the restriction functor $i^{+}$ on the corresponding categories of $\mathcal{D}$-modules is exact, and commutes with pushforwards from open affine subvarieties [Mil, Remark 3.1].
@@
Our strategy for gaining intuition about the $\mathcal{D}_{\widetilde{X}}$-modules arising in the construction of the Jantzen filtration is to illustrate the $\mathfrak{g}$-module structure on their global sections. This will give us an algebraic snapshot as to what is happening at each step in the construction sketched in Section 1.3. The first step is to differentiate the actions (2.7) and (2.8) to obtain a map $\mathcal{U}(\mathfrak{g})\otimes_{\mathcal{Z}(\mathfrak{g})}\mathcal{U}(\mathfrak{h})\rightarrow\Gamma(\widetilde{X},\mathcal{D}_{\widetilde{X}})$. This map provides the $\mathfrak{g}$-module structure on the global sections of $\mathcal{D}_{\widetilde{X}}$-modules. We dedicate this section to the computation of this map.
@@ 2.2. Monodromic $\mathcal{D}_{X}$-modules
An $H$-monodromic $\mathcal{D}_{X}$-module is a weakly $H$-equivariant [^7] A weakly $H$-equivariant $\mathcal{D}_{\widetilde{X}}$-module is an $H$-equivariant sheaf $\mathcal{V}$ equipped with a $\mathcal{D}_{\widetilde{X}}$-module structure so that the isomorphism $\mathrm{act}^{*}\mathcal{V}\rightarrow p^{*}\mathcal{V}$ given by the equivariant sheaf structure on $\mathcal{V}$ is a morphism of $\mathcal{D}_{\widetilde{X}}\boxtimes\mathcal{O}_{H}$-modules. Here $\mathrm{act}:\widetilde{X}\times H\rightarrow\widetilde{X}$ is the action map and $p:\widetilde{X}\times H\rightarrow\widetilde{X}$ is the projection map. For a reference on weakly equivariant $\mathcal{D}$-modules, see [MP98, §4]. $\mathcal{D}_{\widetilde{X}}$-module.
There is an equivalent characterization of $H$-monodromic $\mathcal{D}_{X}$-modules in terms of $H$-invariant differential operators which is established in [BB93, §2.5.2]. This perspective makes the structures of our examples more transparent, so we will take this approach to monodromicity. Below we describe the construction for $\mathfrak{g}=\mathfrak{sl}_{2}(\mathbb{C})$.
@@ 2.3. Verma modules and dual Verma modules
Using the map (2.18) constructed in Section 2.1, we can describe the $\widetilde{\mathcal{U}}$-module structure on various classes of $\mathcal{D}_{\widetilde{X}}$-modules. We will start by examining the $\mathcal{D}_{\widetilde{X}}$-modules $j_{+}\mathcal{O}_{U}$ and $j_{!}\mathcal{O}_{U}$, where $j:U\hookrightarrow\widetilde{X}$ is inclusion of the open union of $N$-orbits
@@
Here the $+$ and !!! indicate the $\mathcal{D}$-module push-forward functors, see [Mil]. These are the $\mathcal{D}_{\widetilde{X}}$-modules which will eventually be endowed with geometric Jantzen filtrations in Section 2.5. In this section, we describe the $\widetilde{\mathcal{U}}$-module structure on $\Gamma(\widetilde{X},j_{+}\mathcal{O}_{U})$ and $\Gamma(\widetilde{X},j_{!}\mathcal{O}_{U})$.
@@
This is a well-defined functor from the category of holonomic $\mathcal{D}_{Y}$-modules to itself [HTT08, Corollary 2.6.8].
@@
    The sheaf $\pi_{*}j_{!}\mathcal{O}_{U}$ is an $H$-monodromic $\mathcal{D}_{X}$-module. For each positive integer $n$, $\pi_{*}j_{!}\mathcal{O}_{U}$ has a subsheaf $(\pi_{*}j_{!}\mathcal{O}_{U})_{n}$ on which $R_{h}$ acts locally by the eigenvalue $n$. These subsheaves are modules over the TDO $\mathcal{D}_{n}$ (Remark 2.4). These are exactly the $\mathcal{D}_{n}$-modules appearing in [Rom21, §6, Fig. 3].
@@ 2.4. The maximal extension $\Xi_{\rho}\mathcal{O}_{U}$
This functor (defined in (2.55) below) extends $j_{+}$ and $j_{!}$ (see (2.57) - (2.58)), so it is a natural way to study both modules $j_{!}\mathcal{O}_{U}$ and $j_{+}\mathcal{O}_{U}$ at once. In this section, we give the construction of $\Xi_{\rho}$, then describe the $\widetilde{\mathcal{U}}$-module structure on $\Gamma(\widetilde{X},\Xi_{\rho}\mathcal{O}_{U})$.
@@
to be the $\mathcal{D}_{U}$-module obtained by twisting $\mathcal{M}_{U}$ by $I^{(n)}$. Note that $f^{s}\mathcal{M}_{U}^{(1)}=\mathcal{M}_{U}$, and that both $I^{(n)}$ and $f^{s}\mathcal{M}_{U}^{(n)}$ have a natural action by $s\in\mathbb{C}[s]/s^{n}$.
@@
Now, we apply this general construction in the setting of our example. Let $\widetilde{X}$ and $U$ be as above (see (2.5) and (2.34)), and let $f_{\rho}$ be the function [^9] This choice of function corresponds to the deformation direction $\rho\in\mathfrak{h}^{*}$, see Remarks 1.3 and 2.9..
@@ 2.4.2. Step 2: $+$-pushforward
annihilates any monomial basis element. (Here $\gamma_{\mathrm{HC}}$ is the Harish-Chandra projection in (2.3).) Hence, the global sections of the submodules of $j_{+}f^{s}\mathcal{O}_{U}^{(n)}$ spanned by monomials $\{x_{1}^{k}x_{2}^{\ell}s^{m}t^{s}\}$ for fixed integers $k+\ell$ have generalised, but not strict, infinitesimal character.